According to last year’s report, a filipino household spends an average of P400/day for food. Suppose you took a sample of 25 households. You determined how much each household spent for the food daily and the results revealed a mean of P380 and a standard deviation of P21.50. with 99% confidence what would be your conclusion



1
Expert's answer
2021-06-08T10:16:59-0400

One sample t-est.

"H_0:\mu=400."

"H_a:\mu\ne400."

Degrees of freedom: "df=25-1=24."

Test statistic: "t=\frac{380-400}{\frac{21.50}{\sqrt{25}}}=-4.65."

P-value: "p=P(T<-4.65)=0.0001."

Since the P-value is less than 0.01, reject the null hypothesis.